2.5points

eBook

Print

References

Check my workCheck My Work button is now enabled

Item 2

Item 2 2.5 points

All of the accounts of the Grass is Greener Company have been adjusted as of December 31, 2018, with the exception of income taxes incurred but not yet recorded. Those account balances appear below. All have normal balances. The estimated income tax rate for the company is 40%.

Cash $ 369,340
Accounts Receivable 789,950
Interest Receivable 5,200
Prepaid Insurance 7,700
Prepaid Rent 13,000
Supplies 225,900
Equipment 690,500
Accumulated Depreciation 143,300
Accounts Payable 299,700
Deferred Revenue 92,600
Income Tax Payable 0
Salaries and Wages Payable 28,800
Notes Payable (long-term) 374,040
Long–Term Debt 247,600
Common Stock 402,200
Retained Earnings 221,800
Dividends 23,600
Service Revenue 958,000
Interest Revenue 132,100
Supplies Expense 354,200
Repairs and Maintenance Expense 262,300
Depreciation Expense 61,350
Rent Expense 32,300
Income Tax Expense Unknown

Required:

  1. Calculate the income before income tax.
  2. Calculate the income tax expense.
  3. Calculate the net income.
Answer #1

a) Income before income tax

Service revenue 958000
Interest revenue 132100
Supplies expense -354200
Repairs and maintenance expense -262300
Depreciation expense -61350
Rent expense -32300
Income before income tax 379950

b) Income tax expense = 379950*40% = 151980

c) Net income = 379950-151980 = 227970